OpenStudy (anonymous):

Simplify the following expression, and rewrite it in an equivalent form with positive exponents. (30x^3y^-5/150x^-4y^2)^2

OpenStudy (anonymous):

x^14\/(25 y^14)

OpenStudy (mertsj):

\[(\frac{30x^3y ^{-5}}{150x ^{-4}y ^{2}})^{2}=(\frac{30x ^{3}x ^{4}}{150y ^{2}y ^{5}})^{2}\]

OpenStudy (mertsj):

\[(\frac{x ^{7}}{5y ^{7}})^{2}=\frac{x ^{14}}{25y ^{14}}\]